Building A is 338 meters taller than building B. If the height of building A is subtracted from twice the height of building​ B, the result is 131 meters. How tall is each​ skyscraper?
During his tennis career in singles​ play, John won 33 fewer tournament A titles than tournament B titles and 22 more tournament C titles than tournament B titles. If he won 20 of these titles​ total, how many times did he win each​ one?
John won
nothing tournament A​ title(s).
John won
nothing tournament B​ title(s).
John won
nothing tournament C​ title(s).

Answer by Theo(13342)

first question:

Building A is 338 meters taller than building B. If the height of building A is subtracted from twice the height of building​ B, the result is 131 meters. How tall is each​ skyscraper?

let a = the height of building A and b = the height of building B.

you get a = b + 338

if you subtract the height of building A from twice the height of building B, then you get:

2b - a = 131.

you have 2 equations to be solved simultaneously.

they are:

a = b + 338
2b - a = 131

in the second equation, replace a with its equivalent value from the first equation to get:

2b - (b + 338) = 131

simplify to get 2b - b - 338 = 131

combine like terms to get b - 338 = 131

add 338 to both sides of the equaiton to get b = 469.

that's the height of building B.

since building A is 338 meters higher than that, then building A must be 807 meters high.

you get building A is 807 meters high and building B is 469 meters high.

building A is 807 - 469 = 338 meters higher than building B.

if you subtract the height of building A from twice the height of building B, you get 2 * 469 - 807 = 938 - 807 = 131 meters difference.

solution to the first question looks good.

building A is 807 meters high and building B is 469 meters high.

second question:

During his tennis career in singles​ play, John won 33 fewer tournament A titles than tournament B titles and 22 more tournament C titles than tournament B titles. If he won 20 of these titles​ total, how many times did he win each​ one?

let a = number of tournament A titles won.
let b = number of tournament B titles won.
let c = number of tournament C titles won.

since total number tournament titles won is 20, you get:

a + b + c = 20

since he won 33 fewer tournament A titles than tourament B titles, you get:

a = b - 33

since he won 22 more tournament C titles than tournament B titles, you get:

c = b + 22.

replacing a and c with their equivalent values from the first 2 equations, you get the third equation becoming:

b - 33 + b + b + 22 = 20

combine like terms to get 3b - 11 = 20

add 11 to both sides to get 3b = 31

divide both sides by 3 to get b = 10 and 1/3.

since a = b - 33, this makes a = - (22 and 2/3).

since c = b + 22, this makes c = 32 and 1/3.

a + b + c = 20, but .....

a is negative which can't be, indicating there is something wrong with the equation.

i suspect that the total number of tournaments won have to be a number greater than or equal to 33 in order for a to be greater than or equal to 0.

i don't believe there is a valid solution to this problem the way it is presented.

first of all, the value of b was not an integer.

second of all, the value of a is negative.

these are two indications that there is something wrong with the equation the way it is presented.

i suspect the problem is in the number of tournaments won, but that's just a guess.